# Distributed Optimization of Clique-Wise Coupled Problems via Three-Operator Splitting

### v2

@@ 1 Introduction
To handle wider applications that involve complex coupling beyond edges, we address a more generic class of distributed optimization—clique-wise coupled optimization problems. This class has been introduced in our recent works [30, 31] with an emphasis on its generalization aspect. In this note, we elucidate additional benefits of this class of problems for performance enhancement and analysis via a new algorithm based on a three-operator splitting [13]. This class of problems is formulated as follows: Consider a multi-agent system with $n$ agents over a time-invariant undirected graph $\mathcal{G}=(\mathcal{N},\mathcal{E})$ with $\mathcal{N}=\{1,\ldots,n\}$ and an edge set $\mathcal{E}$. Let $x_{i}\in\mathbb{R}^{d_{i}}$ represent the $d_{i}$ dimensional decision variable of agent $i$. Then, the following is called a clique-wise coupled optimization problem:
@@
where for all $l\in\mathcal{Q}_{\mathcal{G}}$, $f_{l}:\mathbb{R}^{\sum_{j\in{\mathcal{C}_{l}}}d_{j}}\to\mathbb{R}$ is $L_{l}$-smooth and convex, and $g_{l}:\mathbb{R}^{\sum_{j\in{\mathcal{C}_{l}}}d_{j}}\to\mathbb{R}$ is proper, closed, and convex. For all $i\in\mathcal{N}$, $\hat{f}_{i}:\mathbb{R}^{d_{i}}\to\mathbb{R}$ is $\hat{L}_{i}$-smooth and convex, and $\hat{g}_{i}:\mathbb{R}^{{d_{i}}}\to\mathbb{R}$ is proper, closed, and convex. For the set ${\mathcal{C}_{l}}=\{j_{1},\ldots,j_{|{\mathcal{C}_{l}}|}\}\subset\mathcal{N}$, let $x_{{\mathcal{C}_{l}}}$ denote $x_{{\mathcal{C}_{l}}}=[x_{j_{1}}^{\top},\ldots,x_{j_{|{\mathcal{C}_{l}}|}}^{\top}]^{\top}$. Here, the set ${\mathcal{C}_{l}}$ represents a clique, i.e., a complete subgraph in $\mathcal{G}$ [7]. The set $\mathcal{Q}_{\mathcal{G}}\neq\emptyset$ is a subset of ${\mathcal{Q}_{\mathcal{G}}^{\mathrm{all}}}$, where ${\mathcal{Q}_{\mathcal{G}}^{\mathrm{all}}}$ is the index set of all the cliques in $\mathcal{G}$. For example, in the undirected graph in Fig. 1, we have ${\mathcal{Q}_{\mathcal{G}}^{\mathrm{all}}}=\{1,\ldots,9\}$ and $\mathcal{C}_{1},\ldots,\mathcal{C}_{9}$.
@@ 3.2 Useful properties
    $\mathrm{prox}_{\alpha G}(\mathbf{y})=\mathbf{D}\begin{bmatrix}\mathrm{prox}_{\frac{\alpha}{|\mathcal{Q}_{\mathcal{G}}^{1}|}\hat{g}_{1}}(E_{1}(\mathbf{D}^{\top}\mathbf{D})^{-1}\mathbf{D}^{\top}\mathbf{y}))\\ \vdots\\ \mathrm{prox}_{\frac{\alpha}{|\mathcal{Q}_{\mathcal{G}}^{n}|}\hat{g}_{n}}(E_{n}(\mathbf{D}^{\top}\mathbf{D})^{-1}\mathbf{D}^{\top}\mathbf{y}))\end{bmatrix}$ (15)
@@
    $\mathrm{prox}_{\alpha G}^{\mathbf{Q}}(\mathbf{y})=\mathbf{D}\begin{bmatrix}\mathrm{prox}_{\alpha\hat{g}_{1}}(E_{1}(\mathbf{D}^{\top}\mathbf{D})^{-1}\mathbf{D}^{\top}\mathbf{y}))\\ \vdots\\ \mathrm{prox}_{\alpha\hat{g}_{n}}(E_{n}(\mathbf{D}^{\top}\mathbf{D})^{-1}\mathbf{D}^{\top}\mathbf{y}))\end{bmatrix}$ (16)
@@
    $\frac{\partial}{\partial\mathbf{y}}\sum_{i=1}^{n}\hat{f}_{i}(E_{i}(\mathbf{D}^{\top}\mathbf{D})^{-1}\mathbf{D}^{\top}\mathbf{y})$
    $=\mathbf{D}(\mathbf{D}^{\top}\mathbf{D})^{-1}\begin{bmatrix}\nabla_{x_{1}}\hat{f}_{1}(E_{1}(\mathbf{D}^{\top}\mathbf{D})^{-1}\mathbf{D}^{\top}\mathbf{y})\\ \vdots\\ \nabla_{x_{n}}\hat{f}_{n}(E_{n}(\mathbf{D}^{\top}\mathbf{D})^{-1}\mathbf{D}^{\top}\mathbf{y})\\ \end{bmatrix}$ (17)
@@ 4.1 Algorithm description
$\!\!\!\!\!\!\!\!\!\begin{array}[]{lll}&\mathbf{x}^{k}=\mathrm{prox}_{\sum_{i=1}^{n}\frac{\alpha}{|\mathcal{Q}_{\mathcal{G}}^{i}|}\hat{g}_{i}(\cdot)}(\mathbf{D}^{\top}\mathbf{D})^{-1}\mathbf{D}^{\top}\mathbf{z}^{k}\\ &\mathbf{y}^{k+1/2}=\mathbf{D}\mathbf{x}^{k}\\ &\mathbf{y}^{k+1}=\mathrm{prox}_{\alpha g}(2\mathbf{y}^{k+1/2}-\mathbf{z}^{k}\\ &\quad\quad\quad-\alpha\nabla_{\mathbf{y}}f(\mathbf{y}^{k+1/2})-\alpha\mathbf{D}(\mathbf{D}^{\top}\mathbf{D})^{-1}\nabla_{\mathbf{x}}\hat{f}(\mathbf{x}^{k}))\\ &\mathbf{z}^{k+1}=\mathbf{z}^{k}+\mathbf{y}^{k+1}-\mathbf{y}^{k+1/2},\end{array}$ (28)
@@
For Problem (31), Proposition 1 tells us that the function $\sum_{i=1}^{n}\hat{g}_{i}(E_{i}(\mathbf{D}^{\top}\mathbf{D})^{-1}\mathbf{D}^{\top}\mathbf{y})+\delta_{\mathrm{Im}(\mathbf{D})}(\mathbf{y})$ is proximable for proximable $\hat{g}_{i}$, and the proximal operator can be computed in a distributed fashion. Accordingly, we can directly apply DYS in (9) with $M=I$ to (31). From Proposition 1, setting $x_{i}^{k}=\mathrm{prox}_{\frac{\alpha}{|\mathcal{Q}_{\mathcal{G}}^{i}|}\hat{g}_{i}}(E_{i}(\mathbf{D}^{\top}\mathbf{D})^{-1}\mathbf{D}^{\top}\mathbf{z}^{k})$ gives the distributed algorithm in Algorithm 1 (or (28)).
@@ 4.2 Variable metric version
where we use Propositions 1b and 2. It will turn out in Section 5 that this algorithm shows an interesting connection to other methods as Fig. 3 through ${\boldsymbol{\Phi}}$ in (18). By Lemma 1, a sufficient condition for the convergence is $\alpha\in(0,2/({\max_{l\in\mathcal{Q}_{\mathcal{G}}}\max_{j\in{\mathcal{C}_{l}}}(|\mathcal{Q}_{\mathcal{G}}^{j}|L_{l})+\max_{i\in\mathcal{N}}\hat{L}_{i}}))$.
@@ 5 Revisit of Consensus Optimization as A Clique-Wise Coupled Problem
Figure 3: The relationships among the proposed methods and existing methods.
@@ 5.1 CD-DYS as generalized NIDS
$\begin{array}[]{lll}&\mathbf{x}^{k}=&\mathrm{prox}_{\alpha\hat{g}}(\mathbf{w}^{k})\\ &\mathbf{w}^{k+1}=&\mathbf{w}^{k}-\mathbf{x}^{k}+\widetilde{\mathbf{W}}(2\mathbf{x}^{k}-\mathbf{x}^{k-1}+\alpha\nabla_{\mathbf{x}}\hat{f}(\mathbf{x}^{k-1})-\alpha\nabla_{\mathbf{x}}\hat{f}(\mathbf{x}^{k}))\end{array}$ (42)
@@
$w_{i}^{k+1}=w_{i}^{k}-x_{i}^{k}$
$+\frac{1}{|\mathcal{Q}_{\mathcal{G}}^{i}|}\sum_{l\in\mathcal{Q}_{\mathcal{G}}^{i}}\frac{\mathbf{1}_{|{\mathcal{C}_{l}}|}^{\top}Q_{l}}{\mathbf{1}_{|{\mathcal{C}_{l}}|}^{\top}Q_{l}\mathbf{1}_{|{\mathcal{C}_{l}}|}}(2x^{k}_{\mathcal{C}_{l}}-z_{l}^{k}-\alpha D_{l}\nabla_{\mathbf{x}}\hat{f}(\mathbf{x}^{k})).$
@@
$\mathbf{1}_{|{\mathcal{C}_{l}}|}^{\top}Q_{l}z_{l}^{k+1}=$ $\mathbf{1}_{|{\mathcal{C}_{l}}|}^{\top}Q_{l}(z_{l}^{k}-x_{\mathcal{C}_{l}}^{k})$
$+\mathbf{1}_{|{\mathcal{C}_{l}}|}^{\top}Q_{l}(2x^{k}_{\mathcal{C}_{l}}-z_{l}^{k}-\alpha D_{l}\nabla_{\mathbf{x}}\hat{f}(\mathbf{x}^{k}))$
@@
The original DYS paper [20] states that the NIDS is obtained from the PD3O in [36] (a primal-dual variant of DYS). Meanwhile, in Proposition 5, we rely only on the primal part and obtain $\widetilde{\mathbf{W}}={\boldsymbol{\Phi}}$ as a fixed parameter.
@@ 6 Numerical Experiments
First, we consider a resource allocation problem for the network of $n=20$ agents in [30, Fig. 1] with four communities modeled by cliques and suppose that a resource constraint is imposed on each clique. The clique parameters are given as $\mathcal{Q}_{\mathcal{G}}=\{1,2,3,4\}$ and $\mathcal{C}_{1}=\{1,2,\ldots,6\},\,\mathcal{C}_{2}=\{5,6,\ldots,9\},\,\mathcal{C}_{3}=\{8,9,\ldots,12\},\,\mathcal{C}_{4}=\{9,10,13,14,\ldots,20\}$. Let $x_{i}\geq 0$ for $i\in\mathcal{N}$ be the amount of resources of agent $i$. For $i\in\mathcal{N}$ and $l\in\mathcal{Q}_{\mathcal{G}}$, we set $f_{l}(x_{\mathcal{C}_{l}})=\frac{a_{l}}{2}\|\frac{1}{|{\mathcal{C}_{l}}|}\mathbf{1}_{|{\mathcal{C}_{l}}|}^{\top}x_{\mathcal{C}_{l}}-b_{l}\|^{2}$ with $a_{l}=1$ and $b_{l}$ generated by the uniform distribution with $[0,5]$, $g_{l}(x_{\mathcal{C}_{l}})=\delta_{\mathcal{D}_{l}}(x_{\mathcal{C}_{l}})$ with $\mathcal{D}_{l}=\{x_{\mathcal{C}_{l}}:\sum_{j\in{\mathcal{C}_{l}}}x_{j}=N_{l}\}$, for $(N_{1},\ldots,N_{4})=(5,10,5,15)$, $\hat{f}_{i}(x_{i})=\frac{\hat{a}_{i}}{2}\|x_{i}-\hat{b}_{i}\|^{2}$ with $\hat{a}_{i}=1$ and $\hat{b}_{i}$ generated by the uniform distribution with $[0,1]$, and $\hat{g}_{i}(x_{i})=\delta_{\mathbb{R}_{+}\cup\{0\}}(x_{i})$.
@@
The simulation result is plotted in Fig. 4. The proposed CD-DYS converges much faster than Liang et al. [21] whereas that with $\tau=0.1$ fails to converge. This difference is rooted in the fact that the CD-DYS exploits the community structure of the problem and admits larger stepsizes. Note that to get the largest stepsize for Liang et al., one has to know an upper bound of the norm of dual variables.

### v3

# Yuto Watanabe and Kazunori Sakurama

@@ 1 Introduction
To handle wider applications that involve complex coupling beyond edges, we leverage cliques, complete subgraphs of a graph [9], as a generalization of edges and tackle a more generic class of distributed optimization—clique-wise coupled optimization problems. This class has been introduced in our recent works [10, 11] with an emphasis on its generalization aspect. In this note, we elucidate additional benefits of this class of problems for performance enhancement and analysis via a new algorithm based on a three-operator splitting [12]. This class of problems is formulated as follows: Consider a multi-agent system with $n$ agents over a time-invariant undirected graph $\mathcal{G}=(\mathcal{N},\mathcal{E})$ with $\mathcal{N}=\{1,\ldots,n\}$ and an edge set $\mathcal{E}$. Let $x_{i}\in\mathbb{R}^{d_{i}}$ represent the $d_{i}$ dimensional decision variable of agent $i$. Then, the following is called a clique-wise coupled optimization problem:
@@
where the set ${\mathcal{C}_{l}}\subset\mathcal{N}$ represents a clique, and the set $\mathcal{Q}_{\mathcal{G}}\neq\emptyset$ is a subset of ${\mathcal{Q}_{\mathcal{G}}^{\mathrm{all}}}$, the index set of all the cliques in $\mathcal{G}$. (For example, in the undirected graph in Fig. 1, we have ${\mathcal{Q}_{\mathcal{G}}^{\mathrm{all}}}=\{1,\ldots,9\}$ and $\mathcal{C}_{1},\ldots,\mathcal{C}_{9}$.) For the set ${\mathcal{C}_{l}}=\{j_{1},\ldots,j_{|{\mathcal{C}_{l}}|}\}\subset\mathcal{N}$, let $x_{{\mathcal{C}_{l}}}$ denote $x_{{\mathcal{C}_{l}}}=[x_{j_{1}}^{\top},\ldots,x_{j_{|{\mathcal{C}_{l}}|}}^{\top}]^{\top}$. For all $l\in\mathcal{Q}_{\mathcal{G}}$, $f_{l}:\mathbb{R}^{\sum_{j\in{\mathcal{C}_{l}}}d_{j}}\to\mathbb{R}$ is $L_{l}$-smooth and convex, and $g_{l}:\mathbb{R}^{\sum_{j\in{\mathcal{C}_{l}}}d_{j}}\to\mathbb{R}$ is proper, closed, and convex, where the subscript "$j$" shows the index of agent $j$ in ${\mathcal{C}_{l}}.$ For all $i\in\mathcal{N}$, $\hat{f}_{i}:\mathbb{R}^{d_{i}}\to\mathbb{R}$ is $\hat{L}_{i}$-smooth and convex, and $\hat{g}_{i}:\mathbb{R}^{{d_{i}}}\to\mathbb{R}$ is proper, closed, and convex.
@@ 3.2 Useful properties
    $\mathrm{prox}_{\alpha G}(\mathbf{y})=\mathbf{D}\begin{bmatrix}\mathrm{prox}_{\frac{\alpha}{|\mathcal{Q}_{\mathcal{G}}^{1}|}\hat{g}_{1}}(E_{1}(\mathbf{D}^{\top}\mathbf{D})^{-1}\mathbf{D}^{\top}\mathbf{y}))\\ \vdots\\ \mathrm{prox}_{\frac{\alpha}{|\mathcal{Q}_{\mathcal{G}}^{n}|}\hat{g}_{n}}(E_{n}(\mathbf{D}^{\top}\mathbf{D})^{-1}\mathbf{D}^{\top}\mathbf{y}))\end{bmatrix}.$ (15)
@@
    $\mathrm{prox}_{\alpha G}^{\mathbf{Q}}(\mathbf{y})=\mathbf{D}\begin{bmatrix}\mathrm{prox}_{\alpha\hat{g}_{1}}(E_{1}(\mathbf{D}^{\top}\mathbf{D})^{-1}\mathbf{D}^{\top}\mathbf{y}))\\ \vdots\\ \mathrm{prox}_{\alpha\hat{g}_{n}}(E_{n}(\mathbf{D}^{\top}\mathbf{D})^{-1}\mathbf{D}^{\top}\mathbf{y}))\end{bmatrix}.$ (16)
@@
    $\frac{\partial}{\partial\mathbf{y}}\sum_{i=1}^{n}\hat{f}_{i}(E_{i}(\mathbf{D}^{\top}\mathbf{D})^{-1}\mathbf{D}^{\top}\mathbf{y})=\mathbf{D}(\mathbf{D}^{\top}\mathbf{D})^{-1}\begin{bmatrix}\nabla_{x_{1}}\hat{f}_{1}(E_{1}(\mathbf{D}^{\top}\mathbf{D})^{-1}\mathbf{D}^{\top}\mathbf{y})\\ \vdots\\ \nabla_{x_{n}}\hat{f}_{n}(E_{n}(\mathbf{D}^{\top}\mathbf{D})^{-1}\mathbf{D}^{\top}\mathbf{y})\\ \end{bmatrix}.$ (17)
@@ 4.1 Algorithm description
$\!\!\!\!\!\!\!\!\!\begin{array}[]{lll}&\mathbf{x}^{k}=\mathrm{prox}_{\sum_{i=1}^{n}\frac{\alpha}{|\mathcal{Q}_{\mathcal{G}}^{i}|}\hat{g}_{i}(\cdot)}(\mathbf{D}^{\top}\mathbf{D})^{-1}\mathbf{D}^{\top}\mathbf{z}^{k}\\ &\mathbf{y}^{k+1/2}=\mathbf{D}\mathbf{x}^{k}\\ &\mathbf{y}^{k+1}=\mathrm{prox}_{\alpha g}(2\mathbf{y}^{k+1/2}-\mathbf{z}^{k}-\alpha\nabla_{\mathbf{y}}f(\mathbf{y}^{k+1/2})-\alpha\mathbf{D}(\mathbf{D}^{\top}\mathbf{D})^{-1}\nabla_{\mathbf{x}}\hat{f}(\mathbf{x}^{k}))\\ &\mathbf{z}^{k+1}=\mathbf{z}^{k}+\mathbf{y}^{k+1}-\mathbf{y}^{k+1/2},\end{array}$ (27)
@@
For Problem (30), Proposition 1 tells us that the function $\sum_{i=1}^{n}\hat{g}_{i}(E_{i}(\mathbf{D}^{\top}\mathbf{D})^{-1}\mathbf{D}^{\top}\mathbf{y})+\delta_{\mathrm{Im}(\mathbf{D})}(\mathbf{y})$ is proximable for proximable $\hat{g}_{i}$, and the proximal operator can be computed in a distributed fashion. Accordingly, we can directly apply DYS in (9) with $M=I$ to (30). From Proposition 1, setting $x_{i}^{k}=\mathrm{prox}_{\frac{\alpha}{|\mathcal{Q}_{\mathcal{G}}^{i}|}\hat{g}_{i}}(E_{i}(\mathbf{D}^{\top}\mathbf{D})^{-1}\mathbf{D}^{\top}\mathbf{z}^{k})$ gives the distributed algorithm in Algorithm 1 (or (27)). To implement Algorithm 1, the gradient information $\nabla_{x_{j}}\hat{f}_{j}$ has to be shared in the neighbors. Provided the agents are collaborative within their neighbors, this requirement is not restrictive. We also note that in special cases as consensus optimization, this requirement can be alleviated; see Section 5.
@@ 4.2 Variable metric version
where we have used Propositions 1b and 2. It will turn out in Section 5 that this algorithm shows an interesting connection to other methods as Fig. 3 through ${\boldsymbol{\Phi}}$ in (18). By Lemma 1, a sufficient condition for the convergence is $\alpha\in(0,2/({\max_{l\in\mathcal{Q}_{\mathcal{G}}}\max_{j\in{\mathcal{C}_{l}}}(|\mathcal{Q}_{\mathcal{G}}^{j}|L_{l})+\max_{i\in\mathcal{N}}\hat{L}_{i}}))$.
@@ 5 Revisit of Consensus Optimization as A Clique-Wise Coupled Problem
Figure 3: The relationships between the proposed methods and existing methods.
@@
Note also that the full analysis of Fig. 3 is found in Appendix E.
@@ 5.1 CD-DYS as generalized NIDS
$\begin{array}[]{ll}&\mathbf{x}^{k}=\mathrm{prox}_{\alpha\hat{g}}(\mathbf{w}^{k})\\ &\mathbf{w}^{k+1}=\mathbf{w}^{k}-\mathbf{x}^{k}+\widetilde{\mathbf{W}}(2\mathbf{x}^{k}-\mathbf{x}^{k-1}+\alpha\nabla_{\mathbf{x}}\hat{f}(\mathbf{x}^{k-1})-\alpha\nabla_{\mathbf{x}}\hat{f}(\mathbf{x}^{k}))\end{array}$ (41)
@@
$w_{i}^{k+1}=w_{i}^{k}-x_{i}^{k}+\frac{1}{|\mathcal{Q}_{\mathcal{G}}^{i}|}\sum_{l\in\mathcal{Q}_{\mathcal{G}}^{i}}\frac{\mathbf{1}_{|{\mathcal{C}_{l}}|}^{\top}Q_{l}}{\mathbf{1}_{|{\mathcal{C}_{l}}|}^{\top}Q_{l}\mathbf{1}_{|{\mathcal{C}_{l}}|}}(2x^{k}_{\mathcal{C}_{l}}-z_{l}^{k}-\alpha D_{l}\nabla_{\mathbf{x}}\hat{f}(\mathbf{x}^{k})).$
@@
$\mathbf{1}_{|{\mathcal{C}_{l}}|}^{\top}Q_{l}z_{l}^{k+1}=$ $\mathbf{1}_{|{\mathcal{C}_{l}}|}^{\top}Q_{l}(z_{l}^{k}-x_{\mathcal{C}_{l}}^{k})+\mathbf{1}_{|{\mathcal{C}_{l}}|}^{\top}Q_{l}(2x^{k}_{\mathcal{C}_{l}}-z_{l}^{k}-\alpha D_{l}\nabla_{\mathbf{x}}\hat{f}(\mathbf{x}^{k}))$
@@
The original NIDS paper [4] states that the NIDS is obtained from the PD3O in [29] (a primal-dual variant of DYS). Meanwhile, in Proposition 5, we rely only on the primal part and obtain $\widetilde{\mathbf{W}}={\boldsymbol{\Phi}}$ as a fixed parameter.
@@ 6 Numerical Experiments
First, we consider a resource allocation problem for the network of $n=20$ agents in [10, Fig. 1] with four communities modeled by cliques and, suppose that, for the local consumption, a resource constraint is imposed on each community. The clique parameters are given as $\mathcal{Q}_{\mathcal{G}}=\{1,2,3,4\}$ and $\mathcal{C}_{1}=\{1,2,\ldots,6\},\,\mathcal{C}_{2}=\{5,6,\ldots,9\},\,\mathcal{C}_{3}=\{8,9,\ldots,12\},\,\mathcal{C}_{4}=\{9,10,13,14,\ldots,20\}$. Let $x_{i}\geq 0$ for $i\in\mathcal{N}$ be the amount of resources of agent $i$. For $l\in\mathcal{Q}_{\mathcal{G}}$, we set the clique-wise cost function that can account for the desired resource amount with a weight for each community as $f_{l}(x_{\mathcal{C}_{l}})=\frac{a_{l}}{2}\|\frac{1}{|{\mathcal{C}_{l}}|}\sum_{j\in{\mathcal{C}_{l}}}x_{j}-b_{l}\|^{2}$, with weights $a_{l}=l,\,l\in\{1,\ldots,4\}$ and desired resource amounts $b_{l}$ (generated by the uniform distribution with $[0,50]$), and define the clique-wise resource constraint as $g_{l}(x_{\mathcal{C}_{l}})=\delta_{\mathcal{D}_{l}}(x_{\mathcal{C}_{l}})$ with $\mathcal{D}_{l}=\{x_{\mathcal{C}_{l}}:\sum_{j\in{\mathcal{C}_{l}}}x_{j}\leq N_{l}\}$ for $(N_{1},\ldots,N_{4})=(5,10,5,15)$. For $i\in\mathcal{N}$, we consider the agent’s utility $\hat{f}_{i}(x_{i})=\frac{\hat{a}_{i}}{2}\|x_{i}-\hat{b}_{i}\|^{2}$ with $\hat{a}_{i}=1$, $\hat{b}_{i}$ generated by the uniform distribution with $[0,10]$, and nonnegativity of consumption $\hat{g}_{i}(x_{i})=\delta_{\mathbb{R}_{+}\cup\{0\}}(x_{i})$.
@@
The simulation result is plotted in Fig. 4. The proposed CD-DYS converges much faster than Liang et al. [8] whereas that with $\tau=0.2$ fails to give a descent direction. This difference is rooted in the fact that the CD-DYS exploits the community structure of the problem and admits larger stepsizes. Note that to get the largest stepsize for Liang et al., one has to know an upper bound of the norm of dual variables.